# Majorization for probability distributions, column stochastic matrices and their linear preservers

### v1

@@ 5 Linear operators preserving majorization for the zero sum vectors
A matrix $X\in M_{n,m}$ satisfies condition $(0)$, if $Xa\sim XPa$ for any $a\in\mathbbold{0}^{m}$ and any $P\in P(m)$.
@@
The following lemma shows that we can always find a nonzero $v\in\mathbbold{0}^{n}$ such that $Xv$ has at most two nonzero coordinates. Then the corollary above shows that these nonzero coordinates sum up to zero.
@@
Let $X\in M_{n}$ satisfy $(0)$. Assume that $X^{(i)}\neq X^{(i)}$ for some $i,j\in\mathcal{N}$. Then there is no nonzero $v\in\mathbbold{0}^{n}$ such that $Xv=0$. In particular, $\alpha>0$ in Lemma 5.9.
@@
Similarly, $\alpha+\lambda y_{1}>2\lambda\max\limits_{i}|y_{i}|-\lambda\max\limits_{i}|y_{i}|=\lambda\max\limits_{i}|y_{i}|>\lambda y_{q}$ for any $q\in\mathcal{N}$ and $\alpha+\lambda y_{1}>-\alpha+\lambda y_{2}$. Therefore $\alpha-\lambda y_{1}=\max(b)$.
@@
Case 2. $v_{i}=v_{j}$. Consider arbitrary $\lambda>0$ and $u=v-\lambda(e_{i}-e_{j}),w=v+\lambda(e_{i}-e_{j})$. Observe that $e^{t}u=e^{t}w=e^{t}v=0$. In addition, $\begin{cases}u_{i}=v_{i}-\lambda=v_{j}-\lambda=w_{j};\\ u_{j}=v_{j}+\lambda=v_{i}+\lambda=w_{i};\\ u_{q}=w_{q}=v_{q}\text{, if }q\neq i,j.\end{cases}$ That is, $u=P_{(ij)}w$.
@@ 5.1 Matrices satisfying $(\alpha)$
A matrix $X\in M_{n}$ satisfies condition $(\alpha)$, if there exists $\alpha>0$ such that $X^{(i)}-X^{(j)}\sim\alpha(e_{1}-e_{2})$ for any distinct $i,j\in\mathcal{N}$.
Lemma 5.16 states that if $X\in M_{n}$, $n\geq 3$, satisfies $(0)$ and $X\neq X^{(1)}e^{t}$, then $X$ satisfies $(\alpha)$. In this section, we are going to characterize all square matrices satisfying $(\alpha)$. As we shall see in Theorems 5.22 and 5.25, this property is very similar to preserving majorization for the zero sum vectors.
@@
Assume that $x_{jk}\neq x_{jl},x_{jm}$. Then, by Lemma 5.19, $x_{jl}=x_{jm}$. As $X$ satisfies $(\alpha)$, $x_{ik}\neq x_{il}$ and $x_{jk}\neq x_{jl}$, we obtain that $x_{qk}=x_{ql}$ for any $q\in\mathcal{N}$, $q\neq i,j$. Similarly, $x_{qk}=x_{qm}$ for any $q\in\mathcal{N}$, $q\neq i,j$. It follows that $X^{(l)}=X^{(m)}$ and $X^{(l)}-X^{(m)}=0$, a contradiction. ∎
@@
By Lemma 5.19 in every row of $X$ there are at most two distinct entries. For $n\geq 3$ this immediately proves the required result.
@@
The following theorem characterizes square matrices satisfying $(\alpha)$. In the next section we will show that this essentially characterizes linear operators preserving majorization for the zero sum vectors.
@@
5.2 Characterization of linear operators preserving majorization for the zero sum vectors
@@ 6 Linear operators preserving strong majorization for column stochastic matrices
In this section we give a complete characterization of linear operators preserving strong majorization for column stochastic matrices. We consider a standard decomposition of a linear operator on $M_{n,m}$ into $m^{2}$ linear operators on ${\mathbb{R}}^{n}$.
@@
However, we will discover that every $\Phi^{j}_{i}$ preserves majorization for the zero sum vectors. This will be shown in Section 6.1. Moreover, in the following lemma we show that the operators $\sum\limits_{j=1}^{m}\Phi^{j}_{k}$ do preserve majorization on $\mathbbold{1}^{n}$. Recall that due to Theorem 4.5 this is equivalent to preserving vector majorization in general.
@@ 6.1 Every $\Phi^{j}_{k}$ preserves majorization on $\mathbbold{0}^{n}$
Recall that $a^{+}$ denotes the sum of the positive entries of $a\in{\mathbb{R}}$. That is, $a^{+}=\sum\limits_{i=1}^{n}\max(a_{i},0)$.
@@
Finally, $\phi(b)\sim\phi(Pb)$ for any $b\in\mathbbold{0}^{n}$ and any $P\in P(n)$. By Lemma 6.4 this is equivalent the fact that $\phi$ preserves majorization on $\mathbbold{0}^{n}$. ∎
@@ 6.2 Permutation matrices of $\Phi^{j}_{k}$ coincide
    $n=3$, $\lambda_{1}=\lambda_{2}$ and $P_{1}+P_{2}\leq J$.
@@
If $n=2$, then $P_{2}=J-P_{1}$. It follows that $Y=we^{t}+\lambda_{2}(J-P_{1})=(w+e)e^{t}-\lambda_{2}P_{1}$.
@@
Denote $x=\sum\limits_{q=1}^{m}v^{q}+(\frac{1}{n}\sum\limits_{r\neq i,j}\lambda_{r})e$. If $m=2$, the latter sum is just 00.
@@
$\Phi(A)^{(k)}=\sum\limits_{q=1}^{n}\Phi^{q}_{k}(A^{(q)})=\frac{1}{n}\sum\limits_{r\neq i,j}(v_{r}e^{t}+\lambda_{r}P_{r})e+(v_{i}e^{t}+\lambda_{i}P_{i})e_{l_{1}}+(v_{j}e^{t}+\lambda_{i}P_{j})e_{l_{2}}=$
@@
But $\max(x^{\prime}+2e_{l_{1}})=\max(x^{\prime})+2$, while $\max(x^{\prime}+e_{l_{1}}+e_{l_{2}})\leq\max(x^{\prime})+1$, a contradiction.
@@
Assume that $n\geq 3$. Then by Lemma 6.11 there exist $P^{\prime}\in P(n)$, $v_{1},\ldots,v^{m}\in{\mathbb{R}}^{n}$ and $\lambda_{1},\ldots,\lambda_{m}\in{\mathbb{R}}$ such that
@@
Therefore, $[\Phi^{j}_{k}]=v^{j}e^{t}+\lambda_{j}P,\text{ for any }j\in\mathcal{M}.$ It follows that $\alpha J+\beta Q=(\sum\limits_{j=1}^{m}v^{j})e_{t}+(\sum\limits_{j=1}^{m}\lambda_{j})P$. That is,
$(\alpha e-\sum\limits_{j=1}^{m}v^{j})e_{t}=(\sum\limits_{j=1}^{m}\lambda_{j})P-\beta Q.$
@@ 6.3 Characterization
Let $\Phi(X)=\sum\limits_{j=1}^{m}(e^{t}X^{(j)})S_{j}+PXR\text{ for some }S_{1},\ldots,S_{m}\in M_{n,m},\ R\in M_{m}\text{ and }P\in P(n)$. If $A\in\Omega^{col}_{n,m}$, then $\Phi(A)=\sum\limits_{j=1}^{m}S_{j}+PAR$.
As we shall see next, this condition is not sufficient.
@@
Let $S=\sum\limits_{j=1}^{m}S_{j}$. If $R=0$, then there is nothing to prove. Consider aribitrary $l\in\mathcal{M}$ with $R^{(l)}\neq 0$. Let $k\in\mathcal{M}$ such that $r_{kl}\neq 0$. For arbitrary $g,h\in\mathcal{N}$ and $Q\in P(n)$ define $A=\frac{1}{n}(P^{-1}(e_{g}-e_{h})e_{k}^{t}+J)$ and $B=\frac{1}{n}(P^{-1}Q(e_{g}-e_{h})e_{k}^{t}+J)$. As $A,B\in\Omega^{col}_{n,m}$ and $A\sim^{s}B$, we conclude that $\Phi(A)\sim^{s}\Phi(B)$.
@@
Considering the $j$-th column of (7) we obtain $S^{(l)}+r_{kl}(e_{g}-e_{h})\sim^{s}S^{(l)}+r_{kl}Q(e_{g}-e_{h})$. As $r_{kl}\neq 0$ and $g,h,Q$ are arbitrary, we can apply Lemma 6.14 and conclude that $S^{(l)}=v_{l}e$ for some $v_{l}\in{\mathbb{R}}$.
@@
7 Conclusion and outlook
In Section 3 we have shown that majorization can always be reduced to the column stochastic matrices. Despite that, restriction to column stochastic matrices provides a richer structure of linear preservers, as follows from Section 6. The following example illustrates that.
@@ 7 Conclusion and outlook
On the other hand, consider, for example $A=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ and $B=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}$. Clearly, $A\sim^{s}B$. However, $\Phi(A)=\begin{pmatrix}2&0\\ 0&1\end{pmatrix}$ and $\Phi(B)=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$. It follows that $\Phi$ does not preserve strong majorization, since $\Phi(B)\not\preceq^{s}\Phi(A)$.
@@
We conclude with a list of open problems:
  • Characterization of linear preservers of strong majorization on $\Omega_{n}$.
  • Characterization of linear preservers of other types of majorization on $\Omega^{col}_{n,m}$ and $\Omega_{n}$.
  • Following our observation at the end of Section 5, one could ask the following question:
    What is the meaning of the form
    $\Phi(X)=\sum\limits_{j=1}^{m}(e^{t}X^{(j)})S_{j}+PXR\text{ for some }S_{1},\ldots,S_{m}\in M_{n,m},\ R\in M_{m}\text{ and }P\in P(n),$
    appearing in Theorem 6.16?
    By Corollary 6.18 these operators are not linear preservers of strong majorization on $\Omega^{col}_{n,m}$. What are they?
    Once again, this is the most natural construction that encompasses both types of operators preserving strong majorization, see Theorem 2.10. Similarly to the vector case, these operators preserve strong majorization for the ”zero column-sum matrices”. Indeed, $\Phi(X)=PXR$, whenever $e^{t}X=0^{t}$.
    Does this form completely characterize linear preservers of strong majorization for the ”zero column sum matrices”, or do other preservers exist?
Acknowledgments. The author is grateful to Prof. Alexander Guterman for interesting and valuable discussions on the topic of linear preservers.

### v2

@@ 5 Linear operators preserving majorization for the zero-sum vectors
A matrix $X\in M_{n,m}$ satisfies condition $(0)$ if $Xa\sim XPa$ for any $a\in\mathbbold{0}^{m}$ and any $P\in P(m)$.
@@
The following lemma shows that we can always find a nonzero $v\in\mathbbold{0}^{n}$ such that $Xv$ has at most two nonzero coordinates. Then the corollary above shows that these nonzero coordinates sum to zero.
@@
Let $X\in M_{n}$ satisfy $(0)$. Assume that $X^{(i)}\neq X^{(j)}$ for some $i,j\in\mathcal{N}$. Then there is no nonzero $v\in\mathbbold{0}^{n}$ such that $Xv=0$. In particular, $\alpha>0$ in Lemma 5.9.
@@
Similarly, $\alpha+\lambda y_{1}>2\lambda\max\limits_{i}|y_{i}|-\lambda\max\limits_{i}|y_{i}|=\lambda\max\limits_{i}|y_{i}|>\lambda y_{q}$ for any $q\in\mathcal{N}$ and $\alpha+\lambda y_{1}>-\alpha+\lambda y_{2}$. Therefore $\alpha+\lambda y_{1}=\max(b)$.
@@
Case 2. $v_{i}=v_{j}$. Consider arbitrary $\lambda>0$ and $u=v-\lambda(e_{i}-e_{j}),w=v+\lambda(e_{i}-e_{j})$. Observe that $e^{t}u=e^{t}w=e^{t}v=0$. In addition, $\begin{cases}u_{i}=v_{i}-\lambda=v_{j}-\lambda=w_{j};\\ u_{j}=v_{j}+\lambda=v_{i}+\lambda=w_{i};\\ u_{q}=w_{q}=v_{q}\text{ if }q\neq i,j.\end{cases}$ That is, $u=P_{(ij)}w$.
@@ 5.1 Matrices satisfying $(\alpha)$
A matrix $X\in M_{n}$ satisfies condition $(\alpha)$ if there exists $\alpha>0$ such that $X^{(i)}-X^{(j)}\sim\alpha(e_{1}-e_{2})$ for any distinct $i,j\in\mathcal{N}$.
Lemma 5.16 states that if $X\in M_{n}$, $n\geq 3$, satisfies $(0)$ and $X\neq X^{(1)}e^{t}$, then $X$ satisfies $(\alpha)$. In this section, we are going to characterize all square matrices satisfying $(\alpha)$. As we shall see in Theorems 5.22 and 5.25, this property is very similar to preserving majorization for the zero-sum vectors.
@@
Assume that $x_{jk}\neq x_{jl},x_{jm}$. Then, by Lemma 5.19, $x_{jl}=x_{jm}$. As $x_{ik}\neq x_{il}$, $x_{jk}\neq x_{jl}$ and $X$ satisfies $(\alpha)$, we obtain that $x_{qk}=x_{ql}$ for any $q\in\mathcal{N}$, $q\neq i,j$. Similarly, $x_{qk}=x_{qm}$ for any $q\in\mathcal{N}$, $q\neq i,j$. It follows that $X^{(l)}=X^{(m)}$ and $X^{(l)}-X^{(m)}=0$, a contradiction. ∎
@@
By Lemma 5.19 in every row of $X$ there are at most two distinct entries. For $n\leq 3$ this immediately proves the required result.
@@
The following theorem characterizes square matrices satisfying $(\alpha)$. In the next section we will show that this essentially characterizes linear operators preserving majorization for the zero-sum vectors.
@@
5.2 Characterization of linear operators preserving majorization for the zero-sum vectors
@@ 6 Linear operators preserving strong majorization for column stochastic matrices
In this section, we give a complete characterization of linear operators preserving strong majorization for column stochastic matrices. We consider a standard decomposition of a linear operator on $M_{n,m}$ into $m^{2}$ linear operators on ${\mathbb{R}}^{n}$.
@@
However, we will discover that every $\Phi^{j}_{i}$ preserves majorization for the zero-sum vectors. This will be shown in Section 6.1. Moreover, in the following lemma we show that the operators $\sum\limits_{j=1}^{m}\Phi^{j}_{k}$ do preserve majorization on $\mathbbold{1}^{n}$. Recall that due to Theorem 4.5 this is equivalent to preserving vector majorization in general.
@@ 6.1 Every $\Phi^{j}_{k}$ preserves majorization on $\mathbbold{0}^{n}$
Recall that $a^{+}$ denotes the sum of the positive entries of $a\in{\mathbb{R}}^{n}$. That is, $a^{+}=\sum\limits_{i=1}^{n}\max(a_{i},0)$.
@@
Finally, $\phi(b)\sim\phi(Pb)$ for any $b\in\mathbbold{0}^{n}$ and any $P\in P(n)$. By Lemma 6.4 this is equivalent to the fact that $\phi$ preserves majorization on $\mathbbold{0}^{n}$. ∎
@@ 6.2 Permutation matrices of $\Phi^{j}_{k}$ coincide
    $n=3$, $\lambda_{1}=\lambda_{2}$ and $P_{1}+P_{2}\leq J$ (entry-wise).
@@
If $n=2$, then $P_{2}=J-P_{1}$. It follows that $Y=we^{t}+\lambda_{2}(J-P_{1})=(w+\lambda_{2}e)e^{t}-\lambda_{2}P_{1}$.
@@
Denote $x=\sum\limits_{q=1}^{m}v^{q}+(\frac{1}{n}\sum\limits_{r\neq i,j}\lambda_{r})e$. If $m=2$, then the second summand is just 00.
@@
$\Phi(A)^{(k)}=\sum\limits_{q=1}^{m}\Phi^{q}_{k}(A^{(q)})=\frac{1}{n}\sum\limits_{r\neq i,j}(v^{r}e^{t}+\lambda_{r}P_{r})e+(v^{i}e^{t}+\lambda_{i}P_{i})e_{l_{1}}+(v^{j}e^{t}+\lambda_{i}P_{j})e_{l_{2}}=$
@@
But $\max(x^{\prime}+2e_{l_{1}})=\max(x^{\prime})+2$, while $\max(x^{\prime}+e_{l_{1}}+e_{l_{2}})=\max(x^{\prime})+1$, a contradiction.
@@
Assume that $n\geq 3$. Then by Lemma 6.11 there exist $P^{\prime}\in P(n)$, $v^{1},\ldots,v^{m}\in{\mathbb{R}}^{n}$ and $\lambda_{1},\ldots,\lambda_{m}\in{\mathbb{R}}$ such that
@@
Therefore, $[\Phi^{j}_{k}]=v^{j}e^{t}+\lambda_{j}P,\text{ for any }j\in\mathcal{M}.$ It follows that $\alpha J+\beta Q=(\sum\limits_{j=1}^{m}v^{j})e^{t}+(\sum\limits_{j=1}^{m}\lambda_{j})P$. That is,
$(\alpha e-\sum\limits_{j=1}^{m}v^{j})e^{t}=(\sum\limits_{j=1}^{m}\lambda_{j})P-\beta Q.$
@@ 6.3 Characterization
Let $\Phi(X)=\sum\limits_{j=1}^{m}(e^{t}X^{(j)})S_{j}+PXR\text{ for some }S_{1},\ldots,S_{m}\in M_{n,m},\ R\in M_{m}\text{ and }P\in P(n)$.
If $A\in\Omega^{col}_{n,m}$, then $\Phi(A)=\sum\limits_{j=1}^{m}S_{j}+PAR$.
As we shall see next, the necessary condition provided by Theorem 6.16 is not sufficient.
@@
Let $S=\sum\limits_{j=1}^{m}S_{j}$. If $R=0$, then there is nothing to prove. Consider arbitrary $l\in\mathcal{M}$ with $R^{(l)}\neq 0$. Let $k\in\mathcal{M}$ such that $r_{kl}\neq 0$. For arbitrary $g,h\in\mathcal{N}$ and $Q\in P(n)$ define $A=\frac{1}{n}(P^{-1}(e_{g}-e_{h})e_{k}^{t}+J)$ and $B=\frac{1}{n}(P^{-1}Q(e_{g}-e_{h})e_{k}^{t}+J)$. As $A,B\in\Omega^{col}_{n,m}$ and $A\sim^{s}B$, we conclude that $\Phi(A)\sim^{s}\Phi(B)$.
@@
Considering the $l$-th column of (7) we obtain $S^{(l)}+r_{kl}(e_{g}-e_{h})\sim^{s}S^{(l)}+r_{kl}Q(e_{g}-e_{h})$. As $g,h,Q$ are arbitrary and $r_{kl}\neq 0$, we can apply Lemma 6.14 and conclude that $S^{(l)}=v_{l}e$ for some $v_{l}\in{\mathbb{R}}$.
@@
Note that the form of operators in Theorem 6.16 is a very natural structure that encompasses both types of Li Poon operators in Theorem 2.10. The same is true of the main result: the two forms of operators in Theorem 2.10 are just particular cases of operators preserving majorization on $\Omega^{col}_{n,m}$ from Theorem 6.19.
@@
In Section 3 we have shown that majorization can always be reduced to the column stochastic matrices. Despite that, restriction to column stochastic matrices provides a richer structure of linear preservers, as follows from Theorem 6.19. The following example illustrates that.
@@
On the other hand, consider, for example $A=\begin{pmatrix}1&0\\ 0&0\end{pmatrix}$ and $B=\begin{pmatrix}0&0\\ 1&0\end{pmatrix}$. Clearly, $A\sim^{s}B$. However, $\Phi(A)=\begin{pmatrix}2&0\\ 0&1\end{pmatrix}$ and $\Phi(B)=\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$. It follows that $\Phi$ does not preserve strong majorization, since $\Phi(A)\not\preceq^{s}\Phi(B)$.
@@
Acknowledgments
The author thanks Professor Alexander Guterman for valuable discussions on majorization theory and linear preserver problems.